OpenStudy (anonymous):

Medal . Please Help ? Explain . If 20.0 mL of Ca(OH)2 with an unknown concentration is neutralized by 37.5 mL of 0.124 M HCl, what is the concentration of the Ca(OH)2 solution? Show all of the work needed to solve this problem. Ca(OH)2 + 2 HCl yields 2 H2 O + CaCl2

OpenStudy (anonymous):

1. find n(Ca(OH)2)/n((HCl) (it is 1/2 as you can see from balanced equation) c= n*V n(Ca(OH)2)= 1/2*n(HCl) c(Ca(OH)2)*V(Ca(OH)2)= 1/2*c(HCl)*V(HCl) c(Ca(OH)2)= 1/2*c(HCl)*V(HCl)/V(Ca(OH)2)

OpenStudy (anonymous):

Thank You ! :)